[a] Calculate the percentage of both the axial and equatorial forms of methylcyclohexane. Assume T = 25 °C. [b] Draw the equilibrating chair conformers. [c] Reading from the left to the right, is the equilibrium you’ve just illustrated representative of an exergonic or an endergonic process? [d] Based upon the response to part ‘c’, will ΔG be positive or negative? [e] Highlight the 1,3-diaxial interactions present in axial conformer.


1
Expert's answer
2021-03-18T05:57:29-0400

a)"\%axil=\frac {0.05253}{1+0.05253} ×100=5\%"


"\%Eq=(100-5)=95\%"

b)


c) It is an endogenic process


d) The "\Delta G" will be positive.


e) The axial methyl group creates 7.6 kJ/mol of steric strain due to 1,3-diaxial interactions.